i have tried over and over to figure out this problem, and I get so far then i don't understand it. 4(1/4+1/5)+4
Answer by fcabanski(1391)

PEMDAS is the order of operations. Left to right: Parenthesis, then Exponents, then Multiplication and division, then addition and subtraction.)


4(1/4 + 1/5)+4 has parenthesis, so perform the operation inside first.


1/4 + 1 /5 --- need a common denominator to add fractions. Use 20. For 1/4 20+4 = 5, so multiply the numerator by 5. 1/4 = 5/20.


1/5 = 20 divided by 5 is 4, 1x4=4, so 1/5 = 4/20.


5/20 + 4/20 = 9/20.


4(9/20)+4. Multiplication is next. 4(9/20) is 4*(9/20) is 4/1 * 9/20 (multiply numerator by num and denominator by den) = 36/20.


36/20 + 4 = 36/20 + 4/1


Convert 4/1 into a fraction with 20 in the denominator. 20/1=20 so 4/1 = 4*20/20 = 80/20.


36/20 + 80/20 = 116/20 Simplify (divide top and bottom by 4) 29/5 = 5 and 4/5


You can also complete the problem with decimals. 1/4 = .25 and 1/5 = .2 so it's 4(.25 + .2) + 4 = 4(.45)+4 = 1.8+4 = 5.8

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)